\begin{thebibliography}{}



\bibitem{Barak}
B. Barak, Z. Dvir, A. Wigderson, A. Yehudayoff, Rank bounds for design matrices with applications to combinatorial geometry and locally correctable codes. Proceedings of the 43rd annual ACM symposium on theory of computing, STOC 2011, San Jose, CA, ACM, New York (2011) 519-528.

\bibitem{BDSS}
A. Bhattacharyya, Z. Dvir, A. Shpilka, S. Saraf, tight lower bounds for 2-query lccs over finite fields. Proceedings of FOCS 2011 (2011) 638-647.

\bibitem{BoMo}
P. Borwein, W.O.J Moser, A survey of Sylvester's problem and its generalizations,
{\it{Aequationes Mathematicae}} 40 (1990) 111-135.

\bibitem{BoEd}
W. Bonnice, M. Edelstein, Flats associated with the finite sets in $\mathbb{P}^d$, {\it{Nieuw Archief voor Wiskunde}} 15 (1967) 11-14.

\bibitem{Dvir}
Z. Dvir, S. Saraf, A. Wigderson, Improved rank bounds for design matrices and a new proof of Kelly's theorem. Forum of Mathematics, Sigma (Vol. 2), Cambridge University Press (2014). 

\bibitem{Gallai}
T. Gallai, Solution to Problem 4065, {\it{American Mathematical Monthly}} 51 (1944) 169--171.

\bibitem{GeKr}
J. Geelen and M.E. Kroeker, A Sylvester-Gallai-type theorem for complex-representable matroids. arXiv preprint arXiv:2212.03307 (2022).

\bibitem{Graham}
R.L. Graham, K. Leeb, B.L. Rothschild, Ramsey's theorem for a class of categories, {\it{Advances in Mathematics}}, 8(3) (1972) 417-433.

\bibitem{Hales}
A.W. Hales and R.I. Jewett, Regularity and positional games, {\it{Transactions of the American Mathematical Society}}, 106(2) (1963) 222-229.

\bibitem{Hansen}
S. Hansen, A generalization of a theorem of Sylvester on the lines determined by a finite point set, {\it{Mathematica Scandinavica}} 16 (1965) 175-180.

\bibitem{Hirzebruch}
F. Hirzebruch, Arrangements of lines and algebraic surfaces. In Arithmetic and Geometry (Papers dedicated to I. R. Shafarevich), vol. 2, 113-140, Birkhauser (1983).

\bibitem{Kelly}
L.M. Kelly, A resolution of the Sylvester-Gallai problem of J.-PP. Serre, {\it{Discrete and Computational Geometry}} 1 (1986) 101-104.

\bibitem{Kung}
J.P.S. Kung, The long-line graph of a combinatorial geometry. II. Geometries representable over two fields of different characteristics, {\it{Journal of Combinatorial Theory, Series B}} 50(1) (1990) 41-53.

\bibitem{Murty}
U.S.R. Murty, Matroids with Sylvester property, {\it{Aequationes Mathematicae}} 4 (1970) 44-50.

\bibitem{Sylvester}
J.J. Sylvester, Mathematical Question 11851, {\it{Educational Times}} 59 (1893) 98.







\end{thebibliography}

\documentclass[12pt]{amsart}

\usepackage{graphicx}					
\usepackage{amssymb}
\usepackage{mathtools}
\usepackage{xcolor}
\usepackage{mathabx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{url}
\usepackage{tkz-euclide}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{claim}{Claim}[theorem]

\newtheorem*{sg}{Sylvester-Gallai Theorem}
\newtheorem*{kelly}{Kelly's Theorem}
\newtheorem*{grt}{Geometric Ramsey Theorem}
\newtheorem*{ghjt}{Geometric Hales-Jewett Theorem}

\newcommand{\del}{\setminus}
\newcommand{\con}{/}
\DeclareMathOperator{\si}{si}
\DeclareMathOperator{\cl}{cl}
\DeclareMathOperator{\HJ}{HJ}
\DeclareMathOperator{\R}{R}
\DeclareMathOperator{\AG}{AG}
\DeclareMathOperator{\PG}{PG}
\DeclareMathOperator{\GF}{GF}
\newcommand{\bF}{\mathbb{F}}

\begin{document}

\title{Unavoidable Flats in Matroids Representable over Prime Fields}

\author{Jim Geelen}
\address{Department of Combinatorics and Optimization, University of Waterloo, Waterloo, Canada} \author{Matthew E. Kroeker}
\address{Department of Combinatorics and Optimization, University of Waterloo, Waterloo, Canada} 

\thanks{This research was partially supported by grants from the Office of Naval Research [N00014-10-1-0851] and NSERC [203110-2011] and by an NSERC Postgraduate Scholarship [Application No. PGSD3 - 547508 - 2020]}

\begin{abstract}
We show that, for any prime $p$ and integer $k \geq 2$, a simple $\GF(p)$-representable matroid with sufficiently high rank has a rank-$k$ flat which is either independent in $M$, or is a projective or affine geometry.  As a corollary we obtain a Ramsey-type theorem for $\GF(p)$-representable matroids. For any prime $p$ and integer $k\ge 2$, if we $2$-colour the elements in any 
simple $\GF(p)$-representable matroid with sufficiently high rank, then there is a monochromatic flat with rank $k$.
\end{abstract}

\maketitle

\sloppy

\section{introduction}

In 1944, Gallai proved the following classical result~\cite{Gallai}, originally conjectured by Sylvester fifty years prior~\cite{Sylvester}.

\begin{sg}
Every rank-$3$ real-representable matroid has a two-point line.
\end{sg}

Much attention has been given to the question of how to meaningfully generalize the Sylvester-Gallai Theorem, as well as whether similar phenomena occur in more abstract geometric settings. The following result of Kelly~\cite{Kelly} is a particularly famous example of the latter.

\begin{kelly}
Every rank-$4$ complex-representable matroid has a two-point line.
\end{kelly}

Note that Kelly's rank bound is best-possible: for example, the ternary affine plane is well known to be complex-representable. For more information on the Sylvester-Gallai Theorem and its various extensions and generalizations, see Borwein and Moser's survey~\cite{BoMo}. In this paper, we are  motivated by Sylvester-Gallai-type problems which assert the existence independent flats of rank greater than two. For instance, under what conditions is a matroid guaranteed to contain a three-point plane? For the case of real-representability, Bonnice and Edelstein~\cite{BoEd} deduced the following result as a corollary of Hansen's Theorem~\cite{Hansen}.

\begin{theorem}[\cite{BoEd}, \cite{Hansen}]\label{hansen}
For an integer $k \geq 2$, if $M$ is a simple rank-$(2k-1)$ real-representable matroid, then $M$ contains a rank-$k$ independent flat.
\end{theorem}

The bound on the rank in Theorem~\ref{hansen} is tight, as can be seen by considering the direct sum of $k-1$ lines. This result was extended to complex-representable matroids by Barak, Dvir, Wigderson and Yehudayoff~\cite{Barak} with weaker bounds. Those bounds have since been tightened (see~\cite{Dvir,GeKr}), but there is still room for further improvement.

Projective geometries show that the Sylvester-Gallai theorem does not hold for matroids in general nor even for matroids representable over finite fields. However,  Murty~\cite{Murty} showed that any simple rank-$r$ matroid with fewer than $2^{r}-1$ points has a two-point line, and additionally that the binary projective geometry is the smallest such matroid with no two-point line. A stronger result for matroids representable over prime fields was proved by Bhattacharyya, Dvir, Shpilka, and Saraf~\cite{BDSS} who showed that, for a prime $p$ and sufficiently large integer $r$, any rank-$r$ $\GF(p)$-representable matroid with no two-point line has at least $p^{\Omega(r)}$ points.

In this paper we prove an analogue to Theorem~\ref{hansen} for the class of $\GF(p)$-representable matroids, where $p$ is prime. Our approach is to determine the unavoidable rank-$k$ flats in $\GF(p)$-representable matroids with sufficiently high rank. The following is our main result. 

\begin{theorem}\label{unavoidable}
For every prime $p$ and integer $k \geq 1$ there is an integer $N_{p}(k)$ such that, if $M$ is a simple $\GF(p)$-representable matroid with $r(M) \geq N_{p}(k)$, then $M$ contains a rank-$k$ flat $F$ such that either $F$ is independent in $M$, $M|F \cong \AG(k-1,p)$, or $M|F \cong \PG(k-1,p)$. 
\end{theorem}

Since affine and projective geometries both contain $p$-point lines, if we preclude $p$-point lines, then independent rank-$k$ flats become unavoidable. We can also avoid  affine and projective geometries by considering matroids with girth at least five; note that $\AG(n-1,2)$ has girth four.

\begin{corollary}
For every prime $p$ and integer $k \geq 1$, if $M$ is a simple $\GF(p)$-representable matroid with sufficiently high rank and girth at least five, then $M$ contains a rank-$k$ independent flat.
\end{corollary}

Another consequence of Theorem~\ref{unavoidable} is a Ramsey theorem for $\GF(p)$-representable matroids. 
\begin{corollary}\label{Ramsey}
For a prime $p$ and integer $k\ge 1$, if $M$ is a simple $\GF(p)$-representable matroid with sufficiently large rank then, then for any $2$-colouring of the points of $M$, there is a monochromatic rank-$k$ flat.
\end{corollary}

Note that, by Theorem~\ref{unavoidable}, to prove Corollary~\ref{Ramsey} it suffices to consider the cases where
$M$ is a large independent set or $M$ is a high rank affine or projective geometry.
For independents sets we get a monochromatic flat by majority, and 
for projective and affine geometries we use  
known Ramsey theorems; see Section~3.

\section{Building on Kelly's Proof}

We start by reviewing Kelly's proof~\cite{Kelly} that every rank-$4$ complex-representable matroid has a two-point line. His proof uses a result of Hirzebruch~\cite{Hirzebruch} that every rank-$3$ complex-representable matroid has a line with at most three points. Let $e$ be an element in a simple rank-$4$ complex-representable matroid $M$.  We assume by way of contradiction that $M$ has no $2$-point line. By Hirzebruch's result, 
$M\con e$ contains a line $L$ with exactly  three points, say $p_1$, $p_2$, and $p_3$.
Let $N$ be the restriction of $M$ to the set $L\cup\{e\}$. Then $N$ is a simple rank-$3$ matroid, with no $2$-point line, and $N$ 
comprises three copunctual lines $p_1\cup\{e\}$, $p_2\cup\{e\}$, and $p_3\cup\{e\}$. Kelly's proof is completed by the following result, which is implicit in~\cite{Kelly}. 
\begin{lemma}\label{kelly} For a field $\bF$,
let $N$ be an $\bF$-representable rank-$3$ matroid comprising three co-punctual lines $L_1$, $L_2$, and $L_3$.
If $N$ has no $2$-point line, then $|L_1|=|L_2|=|L_3|$ and $|L_1|-1$ is divisible by the characteristic of $\bF$, and, hence, $\bF$ has positive characteristic.
\end{lemma}

\begin{proof}[Proof sketch.]
Let $e$ be the common point of $L_1$, $L_2$, and $L_3$.
Any two points $e_1\in L_1\setminus\{e\}$ and $e_2\in L_2\setminus \{e\}$ are collinear with a third point $e_3\in L_3\setminus\{e\}$. By fixing a particular choice of element $e_2$ and varying $e_1$ we see that $|L_3|\ge |L_1|$. Then, by symmetry, we see that $L_1$, $L_2$, and $L_3$ all have the same size. 
Fix two distinct elements $a,b\in L_2\setminus\{e\}$. 
Consider the $2$-regular bipartite graph $H$ with bipartition $(L_1,L_3)$ such that $x\in L_1$ is adjacent to $y\in L_3$ if 
either $\{x,a,y\}$ or $\{x,b,y\}$ is a triangle. Thus $H$ is a disjoint union of even cycles. Consider one of these cycles, say $C$.
A straightforward argument considering the representation of $M|(\{a,b\}\cup V(C))$ reveals that the characteristic of $\bF$ is $\frac{1}{2} |V(C)|$.
Since this holds for each component of $H$, we see that the characteristic of $\bF$ divides $|L_1\setminus\{e\}|$, as required.
\end{proof} 

The matroid $N|(\{a,b\}\cup L_1\cup L_3)$, in the proof sketch, is an example of a ``Reid Geometry''. The details regarding the representation of Reid Geometries, which were omitted in our proof sketch, are given explicitly by Kung~\cite[Proposition 2.2]{Kung}. 

In the case that $\bF$ is a finite field of size $p$, where $p$ is prime, the lines $L_1$, $L_2$, and $L_3$, in Lemma~\ref{kelly}, must have length exactly $p+1$, since longer lines are not
GF$(p)$-representable.
\begin{lemma}\label{reid1} For a prime $p$,
let $N$ be a simple $\GF(p)$-representable rank-$3$ matroid comprising three copunctual lines $L_1$, $L_2$, and $L_3$.
If $N$ has no $2$-point line, then the lines $L_1$, $L_2$, and $L_3$ each have exactly $p+1$ points.
\end{lemma}

While Kelly used Reid geometries to find two-point lines, we are interested in independent flats of arbitrary rank. In the next two lemmas, we derive a high-dimensional generalization of Lemma~\ref{reid1}  suitable for our purposes.  Note that throughout this paper, when we refer to a {\it{point}} in a matroid, we mean a rank-$1$ flat (as opposed to an element in the ground set).

\begin{lemma}\label{reid2} For an integer $m\ge 3$ and a prime $p$, 
let $e$ be an element of a  simple $\GF(p)$-representable matroid $M$ such that $\si(M\con e)$ is a connected $m$-element matroid.  If $e$ is not a coloop in $M$ and $M$ has no hyperplane disjoint from $e$ with fewer than $m$ points, then each of the lines of $M$ containing $e$ has length $p+1$. 
\end{lemma}

\begin{proof} 
Note that each hyperplane disjoint from $e$ contains exactly one element from each line containing $e$. Since $e$ is not a coloop, one of those lines, say $L_1$, has length at least $3$. 
Let $a_1$ and $a_2$ be distinct elements in $L_1\setminus \{e\}$ and let $b\in E(M)\setminus L_1$ be chosen arbitrarily.
There is a hyperplane $H$ that contains $a_1$ and $b$ and is disjoint from $e$.
Note that $M|H$ is isomorphic to ${\mathrm{si}}(M/e)$. Since ${\mathrm{si}}(M/e)$ is connected, there is a circuit $C$ of $M|H$ that contains both $a_1$ and $b$.
Let $c\in C\setminus \{a_1,b\}$ and let $L_2$ and $L_3$ be the lines spanned by $\{e,b\}$ and $\{e,c\}$ respectively. 
Let $X=C\setminus \{a_1,b,c\}$ and let $N$ denote the restriction of $M\con X$ to $L_1\cup L_2\cup L_3$. Note that 
$N$ is a simple rank-$3$ matroid and $L_1$, $L_2$, and $L_3$ are lines in $N$ that intersect at $e$. 

Consider any line $L$ of $N$ that does not contain $e$ and let $a$ and $b$ be distinct points of $L$.
Note that $X \cup \{a,b,e\}$ is independent in $M$. Extend $X \cup \{a,b,e\}$ to a basis
$B$ of $M$ and consider the hyperplane $F$ spanned by $B\setminus\{e\}$. 
This hyperplane intersects each of the lines through $e$ in a point, so $F$ contains one element from each of $L_1$, $L_2$, and $L_3$. Therefore $L$ is a $3$-point line in $N$.
Thus every line in $N$ disjoint from $e$  has at least three points. 

The lines spanned by $\{a_1,b\}$ and $\{a_2,b\}$, in $N$, both
intersect $L_3$ in a point distinct from $e$. Therefore $|L_3|\ge 3$ and, by symmetry, we also have $|L_2|\ge 3$. Thus $N$ has no two-point line. Therefore, by Lemma~\ref{reid1}, the lines $L_1$, $L_2$, and $L_3$ each have exactly $p+1$ points. Then, by our choice of $b$, all lines through $e$ in $M$ have length $p+1$.
\end{proof}

We use Lemma \ref{reid2} iteratively to build affine geometries. 

\begin{lemma}\label{kelly2}
For integers $k\ge 2$  and $n \geq t\ge 2$ and a prime $p$, if $M$ is a matroid of rank at least $t+k-1$ such that every rank-$t$ flat in $M$ has at least $n$ points, then either
\begin{itemize}
\item[(i)] $M$ has an $\AG(k-1,p)$-restriction,
\item[(ii)] there is a rank-$(k-1)$ flat $F$ in $M$ such that every rank-$t$ flat in $M\con F$ has at least $n+1$ points, or
\item[(iii)] $M$ has a rank-$t$ flat $F$ such that $M|F$ is not connected.
\end{itemize}
\end{lemma}

\begin{proof}
It suffices to prove the result in the case that $M$ has rank equal to $t+k-1$. Let $C$ be a rank-$(k-1)$ flat in $M$.
Since $M\con C$ has rank $t$, we may assume that
$M\con C$ has exactly $n$ points since otherwise $(ii)$ holds. Moreover, we may assume that $M\con C$ is connected, since otherwise $(iii)$ holds.

\begin{claim} For each $e\in C$ and $f\in E(M)\setminus C$, the line spanned by $\{e,f\}$ has length $p+1$.
\end{claim}

\begin{proof}[Proof of claim.]
There is a rank-$t$ flat $F$ of $M$ such that $f\in F$ and $r(F\cup C)= r(F)+r(C)$. Note that $M|F\cong \si(M\con C)$, so $F$ has $n$ points and $M | F$ is connected.
Let $N$ denote the restriction of $M$ to the flat spanned by $F\cup\{e\}$. Note that $e$ is not a coloop of $N$ since otherwise by considering a hyperplane of $N$ containing $e$ we get a rank-$t$ flat with fewer than $n$ points. Then, by Lemma~\ref{reid2}, every line of $N$ through $e$ has length $p+1$. In particular, the line spanned by $\{e,f\}$ has length $p+1$.
\end{proof}

Let $\{e_1,\ldots,e_{k-1}\} $ be a basis of $C$, let $f\in E(M)\setminus C$, and for each $i\in\{1,\ldots,k-1\}$ let $S_i:=\cl(\{f,e_1,e_2,\ldots,e_i\})\setminus C$.
By the claim, $|S_1|=p$ and, for each $i\in\{1,\ldots,k-2\}$, each point in $S_i$ lifts to $p$ points in $S_{i+1}$. Therefore $|S_{i+1}|=p|S_i|$ and hence $|S_{k-1}|=p^{k-1}$.
It follows that $M|S_{k-1}\cong \AG(k-1,p)$.
\end{proof}

\section{Proof of the Main Theorem}

Lemma~\ref{kelly2} will do most of the work in finding the unavoidable flats in Theorem~\ref{unavoidable}, but two problems still remain to be solved: how do we find a flat with the desired size in the minor; and what do we do with an affine geometry restriction? We resolve both of these issues using Ramsey theory. The following is a consequence of Graham, Leeb and Rothschild's Ramsey theorem for vector spaces over a finite field~\cite{Graham}.

\begin{grt}
For each prime-power $q$ and positive integer $t$ there is an integer $\R_q(t)$ such that , if we two-colour the points in a rank-$\R_q(t)$ projective geometry over
$\GF(q)$, then there will be a monochromatic rank-$t$ flat.
\end{grt}

Our other Ramsey-theoretic tool is a corollary of the Hales-Jewett Theorem~\cite{Hales}.

\begin{ghjt}
For each prime-power $q$ and positive integers $t$ and $k$ there is an integer $\HJ_q(t,k)$ such that , if we $k$-colour the points in a rank-$\HJ_q(t,k)$ affine geometry over
$\GF(q)$, then there will be a monochromatic rank-$t$ flat.
\end{ghjt}

A consequence of this is that coextensions of huge affine geometries contain large affine geometries.

\begin{lemma}\label{affine}
For a prime-power $q$ and integers $k,m\ge 2$ and $n=\HJ_q(k,q^m)$, 
if $J$ is a $m$-element independent set in  $\GF(q)$-representable matroid $M$ and $M/J$ has an $\AG(n-1,q)$-restriction, then $M$ has an $\AG(k-1,q)$-restriction.
\end{lemma}

\begin{proof} We may assume that $M/J$ is isomorphic to $\AG(n-1,q)$.
The matroid $M$ has a $\GF(p)$-representation of the form
$$
A=\bordermatrix{ & J & \cr
 & I & B \cr
& 0& D},
$$ 
where $I$ denotes the $J\times J$ identity matrix and $D$ is a representation of $\AG(n-1,q)$.
We may further assume that the first row of $D$ is the all-ones vector.
Note that $B$ has at most $q^m$ distinct columns, and we assign each element of $M\con J$ a colour according to the associated column of $B$.
By the Geometric Hales-Jewett Theorem, there is a rank-$k$ flat $F$ of $(M/J)$ which is monochromatic under this colouring. By adding linear combinations of the first row of $D$ to the rows of $B$, we may assume that the restriction of $B$ to $F$ is the all-zero matrix. Therefore 
$M|F=(M\con J)|F\cong \AG(k-1,q)$, as desired.
\end{proof}

Lemma~\ref{affine} allows us to refine Lemma~\ref{kelly2}.

\begin{lemma}\label{restriction}
For integers $k \geq 2$, $n \geq t \geq 2$ and a prime $p$, there is an integer $N_{p}'(k,t,n)$ such that, if $M$ is a simple $\GF(p)$-representable matroid with $r(M) \geq N_{p}'(k,t,n)$, then $M$ has either an $\AG(k-1,p)$-restriction or a rank-$t$ flat which is either disconnected or has at most $n$ points.
\end{lemma}

\begin{proof}
Let $t \geq 2$, and observe that, for $n \geq \frac{p^{t}-1}{p-1}$, the integer $N_{p}'(k,t,n)$ exists for all $k \geq 2$. Now let $k \geq 2$, $t \leq n < \frac{p^{t}-1}{p-1}$, and assume that $N_{p}'(i,t,n+1)$ exists for all $i \geq 2$. Let $M$ be a simple $\GF(p)$-representable matroid with $$r(M) \geq t+k-1 + N_{p}'(\HJ_{p}(k,p^{k-1}),t,n+1).$$ We may assume that every rank-$t$ flat in $M$ is connected and has at least $n+1$ points. By Lemma~\ref{kelly2}, there is a rank-$(k-1)$ flat $F$ in $M$ such that every rank-$t$ flat in $M \con F$ has at least $n+2$ points. Moreover, every rank-$t$ flat in $M \con F$ is connected. By induction, $M \con F$ has an $\AG(\HJ_{p}(k,p^{k-1})-1,p)$-restriction. The result now follows by Lemma~\ref{affine}.
\end{proof}

We can now prove Theorem \ref{unavoidable}; to facilitate induction, we prove a slightly stronger result. 
\begin{theorem}
For every prime $p$ and integers $k,t \geq 1$ there is an integer $N_{p}(k,t)$ such that, if $M$ is a simple $\GF(p)$-representable matroid with $r(M) \geq N_{p}(k,t)$, then $M$ contains either a rank-$t$ independent flat or a flat $F$ such that $M|F \cong \AG(k-1,p)$ or $M|F \cong \PG(k-1,p)$.
\end{theorem}

\begin{proof}
For $p$ prime and $k \geq 1$, we prove the result for $N_{p}(k,1)=1$, and, for $t \geq 2$, we recursively define
$$N_{p}(k,t)=N_{p}'(R_{p}(k-1)+1, \, 2N_{p}(k,t-1),\, 2N_{p}(k,t-1)),$$ 
where, by convention, we take $R_{p}(0)=0$. We proceed by induction on $t$, where the base case holds because any point in a matroid forms a rank-$1$ independent flat. Now let $t \geq 2$, and assume that the assertion holds for $N_{p}(k,t-1)$ for every integer $k \geq 1$.

For $k \geq 1$, suppose $M$ is a simple $\GF(p)$-representable matroid with $r(M) \geq N_{p}(k,t)$.
Let $m=2N_{p}(k,t-1)$. Note that, if $F$ is a rank-$m$ flat with exactly $m$ points, then $M|F$ is not connected. Therefore,
applying Lemma~\ref{restriction} gives one of the following two cases.

{\flushleft \textbf{Case 1.} {\em There is a set $A \subseteq E(M)$ such that $M|A \cong \AG(R_{p}(k-1),p)$.}}

Let $F = \cl_{M}(A)$, and let $G \cong \PG(R_{p}(k-1),p)$ be a matroid with $F \subseteq E(G)$ and $G|F = M|F$. Then $E(G) \setminus A$ is a hyperplane in $G$, and so $G \del A \cong \PG(R_{p}(k-1)-1,p)$. By the Geometric Ramsey Theorem, there is a rank-$(k-1)$ flat $H$ in $G \del A$ such that either $H \subseteq E(M)$ or $H \subseteq E(G) \setminus E(M)$. Let $e \in A$ and let $F = E(M) \cap \cl_{G}(H \cup \{e\})$. Then $F$ is a rank-$k$ flat in $M$, and $M|F$ is either an affine or projective geometry.

{\flushleft \textbf{Case 2.} {\em There is a rank-$m$ flat $F$ of $M$ such that $M|F$ is not connected.} }

Let $N$ be a smallest component of $M|F$, and let $F'=F \setminus E(N)$. Then $r_{M}(F') \geq N_{p}(k,t-1)$, and so, by induction, we may assume that $M|F'$ has a rank-$(t-1)$ independent flat $I$. Then for any $e \in E(N)$, $I \cup \{e\}$ is a rank-$t$ independent flat in $M$.
\end{proof}

\end{document}